OpenStudy (anonymous):

can someone help me with this? 3/-5- -2/3

OpenStudy (anonymous):

Should be 1/15. Do you understand how I got that?

OpenStudy (anonymous):

could you explain?

OpenStudy (anonymous):

Yes sure. \[\frac{ 3 }{ -5 }-\frac{ 2 }{ -3 }\] Find the Common denominator. (15)\[\frac{ 9 }{ -15 }- \frac{ 10 }{ -15 }\]Now we multiplied it by 3 and 5 to get to 15 on both sides.\[\frac{ -1 }{ -15 }\]Now we subtracted normally. \[\frac{ 1 }{ 15 }\] The negatives cancel out.

OpenStudy (anonymous):

thank you. can you help on this one? 3/-4+4/9